OpenStudy (anonymous):

Using the difference quotient in the limit definition, find the derivative of (See below)

OpenStudy (anonymous):

\[y=\sqrt{2x-1}\]

myininaya (myininaya):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\] We need to know what f(x+h) is. We are given \[f(x)=\sqrt{2x-1} => f(x+h)=\sqrt{2(x+h)-1}\] so we have \[f'(x)=\lim_{h \rightarrow 0} \frac{\sqrt{2(x+h)-1}-\sqrt{2x-1}}{h}\]

myininaya (myininaya):

now rationalize the numerator by multiplying top and bottom by the conjugate of the top

myininaya (myininaya):

like this: \[f'(x)=\lim_{h \rightarrow 0} \frac{\sqrt{2(x+h)-1}-\sqrt{2x-1}}{h} \cdot \frac{\sqrt{2(x+h)-1}+\sqrt{2x-1}}{\sqrt{2(x+h)-1}+\sqrt{2x-1}}\]

OpenStudy (anonymous):

Correct.

OpenStudy (turingtest):

sorry to but in but how are you with Tayor series myininaya? http://openstudy.com/#/updates/4ecd4838e4b04e045aee647b

myininaya (myininaya):

\[f'(x)= \lim_{h \rightarrow 0}\frac{(2(x+h)-1)-(2x-1)}{h(\sqrt{2(x+h)-1}+\sqrt{2x-1})}\]

myininaya (myininaya):

simplify the top and you will able to cancel something